Calculate the concentration of N2O5 remaining 600 seconds after the start of its decomposition at 65C when its concentration was 0.040 mole/L. The reaction and its rate law are 2 N2O5(g) ® 4 NO2(g) + O2(g) Rate = k[N2O5] with a rate constant of 5.2 x 10-3 s-1
1
Expert's answer
2019-06-18T05:36:48-0400

Solution.

Since the expression for the rate of a chemical reaction is written down the following equation, it means a first order reaction. We write the expression for the rate constant of a chemical reaction of the first order:


"k = \frac{1}{\tau} \times ln{\frac{Co}{C}}"

"C = \frac{Co}{e^{k \times \tau}}"

C = 0.0018 mole/L

C(left) = 0.040 - 0.0018 = 0.0382 mole/L

Answer:

C(left) = 0.0382 mole/L